Fermi National Laboratory


Accelerator Update

Monday August 18
The day shift began with the Tevatron (TeV) in shot setup and with MiniBooNE taking beam.

At 9 AM, the TeV suffered a sector B1 quench. The output of a CAMAC crate power supply sagged, which in turn caused the output of the corrector power supplies to sag.

A Linac expert began tuning on the Linac beam due to energy variances at 9:31 AM. Beam to MiniBooNE had to be turned off until 10:17 AM.

TeV experts began a study period at 11 AM while the Antiproton Source (Pbar) built up its stack.

The TeV suffered a quench during its first ramp at 12 PM. Operators could not find the cause for this quench.

Operations put the TeV through a dry squeeze at 2:08 PM.

TeV studies began at 2:58 PM.

Experts discovered a Booster multiwire, in the 400 MeV line, in the beam. It was pulled out at 4:22 PM.

Operations established a 12x0 TeV study-store at 5:04 PM.

Operations terminated the TeV store and established another 12x0 study-store at 8:48 PM.

Operation put the TeV into shot setup at 9:16 PM and established store 2923 at 11:15 PM. Its initial luminosity was 4.0E31. Pbar resumed stacking.

Tuesday August 19
The midnight shift began with Operations monitoring store 2923, with Pbar stacking, and with MiniBooNE taking beam.

At 4:28 AM, the TeV suffered at sector F4 quench. Two CAMAC crates in the D-zero (D0) experiment's electronics room lost power. Operators replaced the power supply for one of the crates. An expert wanted to come in and investigate the other crate.

Controls experts checked out the D0 crate problems. Operations restored the files for the devices in these crate at 7:35 AM and then ramped the TeV up to 980 GeV.

Operations made D0 ready for access at 9:41 AM. The experimenters were out of access thirty minutes later.

Booster RF stations 5&6 suffered a vacuum burst at 12:10 PM. They both reset okay.

Operation put the TeV into shot setup at 1:16 PM.

At 1:28 PM, the TeV aborted due to a sector F1 kicker (F17K3) not firing. Experts investigated, but could not find a problem. Operations resumed the TeV shot setup at 1:46 PM.

At 3:35 PM, F17K3 pulled the abort again. Experts investigated.

A TeV RF station (TRF8) tripped off at 5:53 PM. This was the third time this station had sparked and tripped off during the day. Experts had been contacted for each trip.

At 6:09 PM, a Booster power supply (ORBMP) tripped off. Operators discovered a LCW leak in the power supply.

Booster leak

Booster experts had the ORBMP LCW leak repaired at 7:30 PM and began drying the supply. The experts discovered more ORBMP problems at 8:12 PM. They had the power supply running normally by 8:48 PM.

D0 experimenters accessed their collision hall at 8:35 PM.

Operations resumed sending beam to MiniBooNE at 8:51 PM.

TRF8 had its power amplifier tube changed and was up and running by 8:53 PM. At 9:08 PM, experts declared the station was operating normally.

D0 came out of access at 10 PM.

Operations put the TeV into shot setup at 11:32 PM.

Wednesday August 20
The midnight shift began with the TeV in shot setup and with MiniBooNE taking beam.

Operations established store 2928 at 1:07 AM with an initial luminosity of 4.2E31. Pbar resumed stacking.

At 1:31 AM, the TeV suffered a sector F4 quench. After much investigation, experts concluded that the quench was related to beam instabilities caused by misalignment of collimators. A separator also sparked at the beginning of the event; this spark was probably the root cause of the quench.

Operations put the TeV through a dry squeeze at 7:45 AM.

Today's Plan
The stack is currently at 89mA. When the stack reaches 100mA, Operations will put the TeV into shot setup.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 8/20/2003   email Fermilab